Solve the resultant and direction of the forces given by component and closed polygon method force 1= 100 grams, due east force 2= 120 grams 30° south of west force 3= 80 grams, 45° South of East?

1
Expert's answer
2022-03-09T10:32:14-0500

"F_x=F_1-F_3\cos30°+F_3\cos45°=52.6,"

"F_y=F_2\sin30°+F_3\sin45°=116.6,"

"F=\sqrt{F_x^2+F_y^2}=127.9,"

"\alpha=\arctan\frac{ F_y }{F_x}=65.7°" South of East.